OpenStudy (anonymous):

An equation is shown below: 8x + 2(x – 7) = 7x + 3x – 14 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation.

OpenStudy (omarbirjas):

8x+2(x-7)=7x+3x-14

OpenStudy (omarbirjas):

ok first lets get rid of the parenthesis by distributing the 2

OpenStudy (omarbirjas):

8x+2x-14=7x+3x-14

OpenStudy (omarbirjas):

now we add all the like terms

OpenStudy (omarbirjas):

10x-14=10x-14

OpenStudy (omarbirjas):

now we isolate the variable x

OpenStudy (omarbirjas):

wait....

OpenStudy (omarbirjas):

how are we gonna solve for x?

OpenStudy (omarbirjas):

the equation is equal on both sides....

OpenStudy (omarbirjas):

ill just end up with 0=-28...which obviously isnt true...

OpenStudy (omarbirjas):

but for part b, it would be the distributive property...

OpenStudy (omarbirjas):

@MilaHutcherson

OpenStudy (anonymous):

thank you @omarbirjas

OpenStudy (omarbirjas):

what...i didnt answer it tho....

OpenStudy (anonymous):

for part b, continue

OpenStudy (omarbirjas):

yea for part a you cant find x, but you can only simplify to 10x-14=10x-14

OpenStudy (anonymous):

@omarbirjas Wait how many solutions are there

OpenStudy (omarbirjas):

none....you cant find x in this equation....

OpenStudy (anonymous):

okay